OpenStudy (anonymous):

What is the rate of change for f(x) = 7 sin x − 1 on the interval from x = 0 to x = pi over 2? @M4thM1nd

OpenStudy (anonymous):

average rate of change?

OpenStudy (anonymous):

f(0) = -1 f(pi/2) = 6

OpenStudy (anonymous):

the average rate of change is: (f(pi/2) - f(0))/(pi/2 - 0) = (6 -(-1))/(pi/2 - 0) = 7/(pi/2) = 14/pi

OpenStudy (anonymous):

Your really good at math @M4thM1nd Help me some more Please?

OpenStudy (anonymous):

hahaha thanks! Sure!

OpenStudy (anonymous):

I'll be back in 10 min

OpenStudy (anonymous):

Alright I tag u in new Q